OpenStudy (anonymous):

integrate sin(x)/sin(4x)

OpenStudy (anonymous):

ok tell me how can i initiate

OpenStudy (lgbasallote):

maybe you can initiate by expanding sin (4x) \[\sin (4x)\implies 2\sin (2x)\cos (2x)\] \[\implies 2(2\sin x \cos x)(\cos^2 x - \sin ^2 x)\] \[\implies 4\sin x \cos x (\cos^2 x - \sin ^2 x)\] does that help?

OpenStudy (anonymous):

yes then in denominator i hav 4cos@cos2@

OpenStudy (anonymous):

ok i got it take sinx as t and then diff it wrt x ok thank u

OpenStudy (lgbasallote):

welcome